Evaluate the limits of


X²-X-2/X(X-2) As X tends to be 2


1
Expert's answer
2021-06-15T12:09:07-0400

Let us evaluate the limits of "\frac{X²-X-2}{X(X-2)}" as "X" tends to 2.


"\lim\limits_{X\to 2}\frac{X²-X-2}{X(X-2)}=\lim\limits_{X\to 2}\frac{(X-2)(X+1)}{X(X-2)}=
\lim\limits_{X\to 2}\frac{X+1}{X}=\frac{3}{2}."